# Analytical coarse grained potential parameterization by Reinforcement Learning for anisotropic cellulose

### v2

@@ Abstract
Cellulose nanocrystals (CNC) are a type of cellulose with excellent mechanical performance and other merit attributes. According to previous reports, hydrogen bonds play a pivotal role in the anisotropic structure of the CNC. Understanding the structure and mechanical behavior of CNC on a mesoscopic scale is critical for the development and manufacture of cellulose materials. However, experimental observations and atomistic simulations are not appropriate on the mesoscopic scale. In this study, we introduce an analytical coarse-grained (CG) potential that is directly parameterized using Reinforcement Learning. Reinforcement Learning (RL) is a power source for industrial and academic applications in many fields. However, its potential has not yet been fully exploited in molecular dynamics. The RL and Boltzmann inversion methods were employed to develop a novel CG model of cellulose to represent its anisotropy and polymer stiffness. The resultant CG model is not limited to the target properties for training, and can present mechanical properties under other circumstances without additional training. This model confirms that RL can construct a CG potential that is both physically explainable and powerful. The training code is available at https://github.com/EiPiFun/rl-cll-cg.
@@ 1.1 Cellulose nanocrystals
As the most abundant natural polymer resource on Earth, cellulose is widely available and eco-friendly, with many natural sources including wood, bamboo, algae, and bacteria[1, 2]. Renewability, low cost, biocompatibility, biodegradability, and good mechanical performance[3, 4] are the outstanding merits of cellulose, making it a good replacement for petroleum-based materials. As a nano-form of cellulose material, cellulose nanocrystals (CNC) can achieve superior performance, such as a high elastic modulus (approximately 200 GPa), low mass density (1.6 $\rm{g/cm^{3}}$), high specific surface area (several hundreds of $\rm{m_{2}/g}$), and a low coefficient of thermal expansion[4]. Owing to their excellent mechanical properties and interactions with other ions and polar molecules[5], CNC have been widely studied and employed in biomedicine[6], water treatment[7], energy storage[8], and humidity sensors[9]. Computational simulations, including ab initio simulations[10], molecular dynamics[11], and finite element method simulations[12], were performed to reveal the molecular details and kinetics of CNC. Experimental[13, 14, 15], computational[16, 17, 12], and analytical[18, 19] approaches have been exploited to understand the mechanical properties of CNC. Hydroxyls and HBonds play pivotal roles in these behaviors[20, 21, 22].
@@
CNC are composed of parallel cellulose chains and their axial elastic modulus is much larger than their transverse modulus. The covalent bonds along the cellulose chains are much stronger than the van der Waals and HBonding (HBonding) interactions in the transverse section. The axial modulus and transverse modulus of CNC are approximately 100-200 GPa and 10-50 GPa, which is confirmed by experiments[4] and simulations[16, 17]. In the transverse section, the CNC are orthotropic and exhibit three characteristic load directions: along, perpendicular to, and slanted to the HBonding planes, as shown in Figure 1. Flat residues and HBonds are the key factors in the transverse anisotropy of cellulose. Wu et al. emphasized the anisotropy of CNC, and they simulated the stretch in three orthogonal directions (along the chain, parallel, and perpendicular to the HBonding planes) at three different strain rates and found that the mechanical properties (elastic modulus and Poisson’s ratio) are highly anisotropic and independent of the strain rate[17]. When a vertical load is applied, the slant CNC model undergoing frictional sliding is ductile, whereas the vertical and horizontal models are fragile. Frictional sliding is an important mechanical behavior of CNC and has been elucidated in previous studies. The friction of the nearby cellulose layers at different velocities, normal loads, and relative angles was also simulated. They found that the number of HBonds and the distances between layers were important factors in this case[23]. Zhang et al. systematically simulated and examined the impact of the loading direction, interfacial moisture, and misalignment on the CNC-CNC interface friction mechanical behavior. The density of HBonds is strongly correlated with interaction energy in their results[24]. The frictional sliding of CNC is critical to their mechanical performance when the loading direction is slanted toward the HBonding planes.
@@
Figure 1: Orthotropic transverse section and characteristic directions. (a) Characteristic directions of orthotropic transverse sections dominated by HBonding planes. (b) Behaviors in characteristic directions. Flat residues and HBonds are critical for cellulose anisotropy. The vertical and horizontal models are dominated by HBonds and van der Waals interactions, resulting in fragile performances. The load of the slant model induced friction and rotation of cellulose residues.
To reflect the performance of CNC on a larger length scale, many coarse-grained (CG) models have been developed with different CG levels for different purposes such as mechanical properties and molecular interactions[25]. Phenomenon such as self-assembly or cellulose nano networks require supra CG models, which means that several cellulose residues become one bead[26, 27, 28] If the mechanical properties of the CNC are considered, one cellulose residue (one residue or two nearby residues in the same chain) is usually mapped to a bead[29, 30]. The directionality of the transverse sections was neglected. Some researchers have emphasized molecular details and interactions; thus, they will preserve more molecular details and map one cellulose residue to more than one bead instead of several residues into one bead[31, 32, 33, 34].
However, the CG models of cellulose mentioned previously could not reveal the HBond interactions and resultant directionality. Several studies have presented directionality using anisotropic particles[26, 35] or additional bonds that are not supported by real molecular structures to compensate for the directionality [36, 37, 38, 39]. Shishehbor et al. developed a CG CNC by matching the axial elastic modulus and transverse stress-strain curve to investigate the interface interaction and failure mechanism[38]. The CG mappings used by Shishehbor et al. neglected the flat structure of cellulose residue. Anisotropic particles and bonded (BD) directionality are unsuitable when a transverse failure mechanism with frictional sliding is considered in the CG model. Nonetheless, presenting the kinetics, energy, and mechanical properties of CG models is challenging[33].
@@ 1.2 Machine Learning parameterization and Reinforcement Learning
The use of Reinforcement Learning (RL) in CG simulations and potential parameterization is still in its early stages[49]. Wu et al. integrated graph attention neural networks and RL to learn CG mapping[50]. Chandra et al. used a continuous action Monte Carlo Tree Search[51] powered by RL to parameterize an analytical hybrid bond-order coarse-grained force field to simulate the aggregation of liquid-liquid mixtures[49]. RL imitates the learning of creatures and is promising for control and optimization[52, 53]. The model-free RL algorithm can deal with strongly nonlinear problems without prior expert knowledge and automatically perform trial-and-error learning procedures. More information on the RL scheme is presented in the methods section.
@@ 2.1 Mapping and topology
The mapping emphasizes the flat residues of the cellulose chains and the HBond interactions. One cellulose residue was mapped to one backbone CL1 bead and two branched CL2 and CL3 beads, as shown in Figure 2,. Two features are considered: the distance between CL2 and CL3 in a residue and the angle of CL2-CL3 and the backbone. The first distance feature should be sufficiently long to bear the torque and present flat residues, whereas the second angle feature should be approximately 90. In the crystalline state, the angles of the nearby CL2:CL3-CL1(blue:orange-green) and CL3:CL2-CL1(orange:blue-green) (colons indicate that these two beads are not bonded together, whereas dash indicates bonded) are close to $180^{\circ}$, as illustrated in Figure 2.
@@
Figure 2: Mapping and topology. (a) Mapping and (b) mapped overlays for the cellulose residue and (c) laminar HBonding plane. One cellulose residue was mapped to one backbone CL1 bead and two branched CL2 and CL3 beads. The basic idea behind this mapping is to preserve the flat structure and HBonds within the HBonding layer. Qualitatively, the distance between CL2 and CL3 from the same residue was sufficient and the branched beads were expected to be perpendicular to the backbones. Under this mapping, a layer of CNC was mapped as a network structure, where the HBond interactions were stressed. (d) BD topology. BD interactions are required to preserve molecular structure and performance under stretching, bending, and twisting loads. (e) NB topology. For HBond (HB) potentials, two symmetric A:H-D relations (CL2:CL3-CL1 and CL3:CL2-CL1) were defined to represent HBonds. HB potentials were determined using both the A:D distance and A:H-D angle.
Based on this mapping, we applied three types of bonds, six types of angles, and three types of improper dihedrals as the BD interactions (Figure 2). The BD interactions were designed to simulate and restrain the deformations under stretching, bending, and twisting loads. To ensure restraint in the aperiodic chain (which was used for the polymer stiffness simulations during training), the CL2-CL1-CL1 and CL3-CL1-CL1 angles were categorized. A type of directional NB interaction was introduced to describe HBonds, which requires the Acceptor, Hydrogen and Donor beads. Two A:H-D topologies were determined based on the importance of HBonding and symmetries: CL2:CL3-CL1 and CL3:CL2-CL1. This directional interaction is affected by the A:H-D angle and A-D distance[54].
@@ 2.2 Potentials
$V_{System}=V_{Bond}+V_{Angle}+V_{Improper}+V_{non_{b}onded}$ (1)
@@
$V_{HB}=\cases{\epsilon_{HB}[5{(\frac{\sigma_{HB}}{r})}^{12}-6{(\frac{\sigma_{HB}}{r})}^{10}]{{\rm cos}^{n}(\theta-\theta_{0}+\pi)}&, $0<r\leq r_{in}\,\&\theta-\theta_{0}+\pi\geq\theta_{c}$\cr S(i)\cdot\epsilon_{HB}[5{(\frac{\sigma_{HB}}{r})}^{12}-6{(\frac{\sigma_{HB}}{r})}^{10}]{{\rm cos}^{n}(\theta-\theta_{0}+\pi)}&, $r_{in}<r\leq r_{out}\&\theta-\theta_{0}+\pi\geq\theta_{c}$\cr}$ (7)
@@
$S(i)=\frac{{(r_{out}^{2}-r^{2})}^{2}[r_{out}^{2}+2r^{2}-3r_{in}^{2}]}{{(r_{out}^{2}-r_{in}^{2})}^{3}}$ (8)
The LJ potential is a 12-6 formula of Lennard-Jones interactions, and the HB potential is a 12-10 formula with scaling by the power of the cosine of the A:H-D angle. Technically, these are LAMMPS[54] pair styles. The LJ potential is the lj/smooth with a shift, whereas the HB potential is a modified version of the hbond/dreiding/lj. lj/smooth is designed to smooth force to zero at the cutoff radius using a polylinear expression; this modified version of hbond/dreiding/lj ensures that potential energy reach minima at $\theta_{0}$ instead of $\pi$. The energy and force curves of the HB and LJ potentials are presented in Figure SP9.
The equilibrium A:H-D angles in the CG mapping were designed to be close to $\pi$. Even though we generalize it by adding an equilibrium angle of A:H-D, an equilibrium angle close to $\pi$ is suggested to stabilize the conical energy surface.

### v3

@@ Abstract
Cellulose nanocrystals (CNCs) are a type of cellulose with excellent mechanical performance and other merit attributes. According to previous reports, hydrogen bonds play a pivotal role in the anisotropic structure of the CNC. Understanding the structure and mechanical behavior of CNC on a mesoscopic scale is critical for the development and manufacture of cellulose materials. However, experimental observations and atomistic simulations are not appropriate on the mesoscopic scale. In this study, we introduce an analytical coarse-grained (CG) potential following an extended bottom-up approach that is directly parameterized using Reinforcement Learning (RL). RL is a powerful tool for industrial and academic applications in various fields. Nevertheless, the potential of RL has not yet been fully exploited in the field of molecular dynamics. The RL and Boltzmann inversion methods were employed to develop a novel CG model of cellulose to represent its anisotropy and polymer stiffness. The resultant CG model is not limited to the target properties for training, and can reproduce the dynamics mechanical properties under other circumstances without additional training. This model confirms that RL can construct a CG potential that is both physically explainable and powerful. The training code is available at https://github.com/EiPiFun/rl-cll-cg.
@@ 1.1 Cellulose nanocrystals
As the most abundant natural polymer resource on Earth, cellulose is widely available and eco-friendly, with numerous natural sources including wood, bamboo, algae, and bacteria[1, 2]. Renewability, low cost, biocompatibility, biodegradability, and good mechanical performance[3, 4] are the outstanding merits of cellulose, making it a promising alternative to petroleum-based materials. As a nano-form of cellulose material, cellulose nanocrystal (CNC) can achieve superior performance, such as high elastic modulus (approximately 200 GPa), low mass density (1.6 $\rm{g/cm^{3}}$), high specific surface area (several hundreds of $\rm{m_{2}/g}$), and low coefficient of thermal expansion[4]. Owing to their excellent mechanical properties and interactions with other ions and polar molecules[5], CNCs have been widely studied and employed in biomedicine[6], water treatment[7], energy storage[8], and humidity sensors[9]. Computational simulations, including ab initio simulations[10], molecular dynamics[11], and finite element method simulations[12], were performed to reveal the molecular details and kinetics of CNC. Experimental[13, 14, 15], computational[16, 17, 12], and analytical[18, 19] approaches have been exploited to understand the mechanical properties of CNC. Hydroxyls and HBonds play pivotal roles in these behaviors[20, 21, 22].
@@
CNCs are composed of parallel cellulose chains and their axial elastic modulus is much larger than their transverse modulus. The covalent bonds along the cellulose chains are much stronger than the van der Waals and hydrogen bonding (HBonding) interactions in the transverse section. The axial modulus and transverse modulus of CNC are approximately 100-200 GPa and 10-50 GPa, which is confirmed by experiments[4] and simulations[16, 17]. In the transverse section, the CNCs are orthotropic and exhibit three characteristic load directions: along, perpendicular to, and slanted to the HBonding planes, as shown in Figure 1. Flat residues and HBonds are the key factors in the transverse anisotropy of cellulose. Wu et al. emphasized the anisotropy of CNC, and they simulated the stretch in three orthogonal directions (along the chain, parallel, and perpendicular to the HBonding planes) at three different strain rates and found that the mechanical properties (elastic modulus and Poisson’s ratio) are highly anisotropic and independent of the strain rate[17]. When a vertical load is applied, the slant CNC model undergoing frictional sliding is ductile, whereas the vertical and horizontal models are fragile. Friction is an important mechanical behavior of CNC and has been elucidated in previous studies. The friction of the nearby cellulose layers at different velocities, normal loads, and relative angles was also simulated. They found that the number of HBonds and the distances between layers were important factors in this case[23]. Zhang et al. systematically simulated and examined the impact of the loading direction, interfacial moisture, and misalignment on the CNC-CNC interface friction mechanical behavior. The density of HBonds is strongly correlated with interaction energy in their results[24]. The frictional sliding of CNC is critical to their mechanical performance when the loading direction is slanted toward the HBonding planes.
@@
Figure 1: Orthotropic transverse section and characteristic directions. (a) Characteristic directions of orthotropic transverse sections dominated by HBonding planes. (b) Behaviors in characteristic directions. Flat residues and HBonds are critical for cellulose anisotropy. The vertical and horizontal models are dominated by HBonds and van der Waals interactions, resulting in brittle failure. The load of the slant model induced friction and rotation of cellulose residues.
To reflect the performance of CNC on a larger length scale, many coarse-grained (CG) models have been developed with different CG levels for different purposes such as mechanical properties and molecular interactions[25]. Phenomenon such as self-assembly or cellulose nano networks require supra CG models, in which several cellulose residues are mapped into a single bead[26, 27, 28] If the mechanical properties of the CNC are considered, one cellulose residue (one residue or two nearby residues in the same chain) is usually mapped to a bead[29, 30]. The directionality of the transverse sections was neglected. Some researchers have emphasized molecular details and interactions; thus, they will preserve more molecular details and map one cellulose residue to more than one bead instead of several residues into one bead[31, 32, 33, 34].
However, the CG models of cellulose mentioned previously could not reveal the HBond interactions and resultant directionality. Several studies have presented directionality using anisotropic particles[26, 35] or additional bonds that do not correspond to real molecular structures to compensate for the directionality [36, 37, 38, 39]. Shishehbor et al. developed a CG CNC by matching the axial elastic modulus and transverse stress-strain curve to investigate the interface interaction and failure mechanism[38]. The CG mappings used by Shishehbor et al. neglected the flat structure of cellulose residue. Anisotropic particles and bonded (BD) directionality are unsuitable when a transverse failure mechanism with frictional sliding is considered in the CG model. Nonetheless, presenting the kinetics, energy, and mechanical properties of CG models is challenging[33].
@@ 1.2 Machine Learning parameterization and Reinforcement Learning
The use of Reinforcement Learning (RL) in CG simulations and potential parameterization is still in its early stages[49]. Wu et al. integrated graph attention neural networks and RL to learn CG mapping[50]. Chandra et al. used a continuous action Monte Carlo Tree Search[51] powered by RL to parameterize an analytical hybrid bond-order CG force field to simulate the aggregation of liquid-liquid mixtures[49]. RL imitates the learning of creatures and is promising for control and optimization[52, 53]. The model-free RL algorithm can deal with strongly nonlinear problems without prior expert knowledge and automatically perform trial-and-error learning procedures. More information on the RL scheme is presented in the methods section.
@@
1.3 Coarse grained molecular dynamics
To better illustrate the characteristics and advantages of this CNC CG models, several classical CG modeling methods were compared as baselines in the following sections, and a brief introduction of CG modeling was necessary. CG models focus on essential features, are a compromise between accuracy/scale and computational efficiency comparing to AA models, such as the CG models of CNC on a large spatial and time scales[54]. Together with the reduction of the degrees of freedom by mapping process (replacing several atoms with a bead) and the often elimination of electrostatics interactions, the smoother and shorter energy surface from averaging over atomic structure allow them to use much longer integration timestep. Reduction + smooth energy surface + longer timestep are not only significantly fast but also beneficial for sampling collectively emergent mesoscopic behaviors or observable macroscopic properties[55]. Thereby, CG models demonstrate the relationship between molecular scale and performances in larger scales. Although “coarse grained”, they are “fine grained” compared to continuum model and provide insights into molecular fluctuations and interactions. Well-designed CG models are especially helpful to understand the essential aspects of physical phenomena. However, the shortcomings also comes from the reduction; dynamics, non-equilibrium processes, consistence, representability, and transferability[56]. Most of the CG models concentrate on equilibrium trajectories. Pressure is an example to these problems: with much lower number of particles and velocities, even with the help of Brownian motions/Langevin dynamics, it is still hard to confirm the pressure consistence considering entropy and enthalpy, especially when transferred to a new environment. The representability problems of CG models especially for thermodynamics is the research front of many scientists[57, 58, 59, 60, 61].
Top-down and bottom-up are the two main paradigms for CG modeling[62]. Roughly, top-down emphasize on the emergent phenomena while bottom-up are on the basis of microscopic origins. Top-down methods are not directly connected to microscopic details, they usually utilize simple analytical functions and are hard to provide quantitative insights. Bottom-up approaches are usually structure-based, aiming at the statistical mechanics properties including many body potential of mean force, mapped probability distributions. To more accurately approximate, bottom-up usually utilize complex functions. Thus bottom-up is capable to provide quantitative description. The distinctions of there two paradigms are increasingly blurred. For instance, as a famous top-down CG potential, MARTINI also incorporate elements of bottom-up strategies[63].
MARTINI successfully reproduced the molecular structure of bilayers, hydrocarbons, and short peptides. Except for top-down MARTINI, bottom-up Iterative Boltzmann Inversion (IBI)[64], Relative Entropy (RE)[65], and Force Matching (FM)[66] were also compared as baselines. IBI is a structure-based correlation function approach to reproduce atomic radial distribution function. In IBI, all interactions are assumed as independent, thus IBI is often appropriate for ”dilute” NB pairs and ”stiff” BD interactions. RE is a variational approach to minimize the Kullback-Leibler divergence of mapped probability distribution (the configuration-dependent entropy) between AA and CG iteratively. FM is a non-iterative variational but focus on mean force of NB pairs and could be directly solved. Thereby, the FM does not guarantee to reproduce atomic distributions More descriptions and comparisons of CG modeling methods could be referred in the wonderful reviews[54, 55, 56, 62].
Although MARTINI and IBI have already been applied in the cellulose[33, 31, 67, 68], prior researches did not concentrate on the anisotropic mechanical properties, which will be further presented in the following sections. This CG model of CNC follow an extended bottom-up approach: focusing on the molecular details and mapped probability distribution. But it use simple analytical potential functions and parameterize them to reproduce mechanical properties, which is quite similar to the top-down performance-oriented philosophy. And the exploration of dynamics in CG scale is challenging. On the perspective of CG modeling, this is a hybrid-hybrid approach: hybrid both in the bottom-up/top-down approach and equilibrium/non-equilibrium contexts.
@@ 2.1 Mapping and topology
The mapping emphasizes the flat residues of the cellulose chains and the HBond interactions. One cellulose residue was mapped to one backbone CL1 bead and two branched CL2 and CL3 beads (Figure 2). Two features are considered: the distance between CL2 and CL3 in a residue and the angle of CL2-CL3 and the backbone. The first distance feature should be sufficiently long to withstand the torque and present flat residues, whereas the second angle feature should be approximately 90. In the crystalline state, the angles of the nearby CL2:CL3-CL1(blue:orange-green) and CL3:CL2-CL1(orange:blue-green) (colons indicate that these two beads are not bonded together, whereas dash indicates bonded) are close to $180^{\circ}$, as illustrated in Figure 2.
@@
Figure 2: Mapping and topology. (a) Mapping and (b) mapped overlays for the cellulose residue and (c) laminar HBonding plane. One cellulose residue was mapped to one backbone CL1 bead and two branched CL2 and CL3 beads. The basic idea behind this mapping is to preserve the flat structure and HBonds within the HBonding layer. Qualitatively, the distance between CL2 and CL3 from the same residue was sufficient and the branched beads were expected to be perpendicular to the backbones. Under this mapping, a layer of CNC was mapped as a network structure, where the HBond interactions were stressed. (d) BD topology. BD interactions are required to preserve molecular structure and performance under stretching, bending, and twisting loads. (e) NB topology. For HBond (HB) potentials, two symmetric A:H-D relations (CL2:CL3-CL1 and CL3:CL2-CL1) were defined to describe HBonds. HB potentials were defined by using both the A:D distance and A:H-D angle.
Based on this mapping, we applied three types of bonds, six types of angles, and three types of improper dihedrals as the BD interactions (Figure 2). The BD interactions were designed to describe and restrain the deformations under stretching, bending, and twisting loads. To ensure restraint in the aperiodic chain (which was used for the polymer stiffness simulations during training), the CL2-CL1-CL1 and CL3-CL1-CL1 angles were categorized. A type of directional NB interaction was introduced to describe HBonds, which requires the Acceptor, Hydrogen and Donor beads. Two A:H-D topologies were determined based on the importance of HBonding and symmetries: CL2:CL3-CL1 and CL3:CL2-CL1. This directional interaction is affected by the A:H-D angle and A-D distance[69].
@@ 2.2 Potentials
$V_{System}=V_{Bond}+V_{Angle}+V_{Improper}+V_{Nonbonded}$ (1)
@@
$V_{HB}=\cases{\epsilon_{HB}[5{(\frac{\sigma_{HB}}{r})}^{12}-6{(\frac{\sigma_{HB}}{r})}^{10}]{{\rm cos}^{n}(\theta-\theta_{0}+\pi)}&, $0<r\leq r_{in}\,\&\,\theta-\theta_{0}+\pi\geq\theta_{c}$\cr S(i)\cdot\epsilon_{HB}[5{(\frac{\sigma_{HB}}{r})}^{12}-6{(\frac{\sigma_{HB}}{r})}^{10}]{{\rm cos}^{n}(\theta-\theta_{0}+\pi)}&, $r_{in}<r\leq r_{c}\,\&\,\theta-\theta_{0}+\pi\geq\theta_{c}$\cr}$ (7)
@@
$S(i)=\frac{{(r_{c}^{2}-r^{2})}^{2}[r_{c}^{2}+2r^{2}-3r_{in}^{2}]}{{(r_{c}^{2}-r_{in}^{2})}^{3}}$ (8)
The LJ potential is a 12-6 formula of Lennard-Jones interactions, and the HB potential is a 12-10 formula with scaling by the power of the cosine of the A:H-D angle. Technically, these are LAMMPS[69] pair styles. The LJ potential is the lj/smooth with a shift, whereas the HB potential is a modified version of the hbond/dreiding/lj with a shift. lj/smooth is designed to smooth force to zero at the cutoff radius using a polylinear expression; this modified version of hbond/dreiding/lj ensures that potential energy reach minima at $\theta_{0}$ instead of $\pi$. It should be emphasized that smoothing part between $r_{in}$ and $r_{cut}$ ensure the continuity of both force and energy. The energy and force curves of the HB and LJ potentials are illustrated in Figure Supplementary Information.
The equilibrium A:H-D angles in the CG mapping were designed to be close to $\pi$. Even though we generalize the HB potential by adding an equilibrium angle of A:H-D, an equilibrium angle close to $\pi$ is suggested to stabilize the conical energy surface.
The reason we use Lennard-Jones in th CG level was only for simplicity and the easiness to stabilize structure due to the single stationary point. The fourth power of cosine in the HB potential was the same as origin paper which proposed this formula first. For the artifacts (cut, smoothing, and directional HB potential), the sensitivity and ablation checks were performed based on trained coefficients represented in the following sections: Half/double energy coefficients (while distance coefficients were kept), 0.8/1.2 times distance coefficients (while energy coefficients were kept), different n power of cosine, and removal of HB potential (Table Supplementary Information, Table Supplementary Information). These sensitivity and ablation tests confirmed the effectiveness of potentials.